# System Architecting for GEO Communication Satellite Considering On-Orbit Refueling

### v2

@@ Abstract
$Rel$  = reliability function of a satellite
@@
$t$  = time in discretized frame ($=T/\Delta T$)
@@ 2 Problem Description
On the one hand, reducing the initial propellant mass leads to a smaller tank compared to traditional designs, thereby decreasing the satellite’s overall structural mass. This reduction in structural mass may relax the performance requirements on the reaction wheels. Overall, reducing initial propellant mass can lower mission costs—encompassing development, manufacturing, and launch. Additionally, spending less on system acquisition helps mitigate financial losses in the event of mission failure.
@@
$\underset{\mathbf{x}}{\text{maximize}}\ \mathbf{J}=\left[E_{V},\ E_{V}/\sigma_{V}\right]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0},}$ (1)
@@
$\mathbf{x}=\left[T_{\text{life}},\ m_{p,\text{des}}\right]\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{X}}.$ (2)
@@ 3.1 Satellite Sizing and Cost Elements
Here, Eqs. (5) and (6) adopt the relations referred to [7]. In particular, Eq. (5) implies $\kappa$increase in mass for each additional year of design lifetime beyond three years. We then model $m_{\text{adcs}}$ and $m_{\text{str}}$ as proportional to the dry mass, as shown in Eq. (7).
@@ 3.2.3 Decision
Decision 2 is triggered after OOR Service is selected in Decision 1. It occurs $t_{\text{oor}}$ ($<t_{\text{rep}}$) time steps before the projected depletion. This event determines the quantity of propellant for refueling. The life-extension period provided by OOR Service is determined as the period with the highest utility, similar to Decision 1. When a specific life-extension duration is selected, OOR Service with the corresponding propellant amount ($m_{p,\text{oor}}$) is planned after $t_{\text{oor}}$ time steps.
@@ 3.2.4 Action
where $\Delta V_{\text{ot}}$ and $\Delta V_{\text{ot},\text{ideal}}$ represent the actual and ideal $\Delta V$ required for orbital transfer from the initial orbit to the target orbit, respectively, and $\Delta V_{\text{err}}$ is the additional $\Delta V$ needed for injection error correction. Here, $\mathcal{H}\left(\sigma_{\text{oi}}\right)$ denotes a half-normal distribution derived by taking the absolute value of a normal distribution with mean 00 and standard deviation $\sigma_{\text{oi}}$.
@@ 4.2 Surrogate Model Formulation
$\overline{V}_{i}=\frac{1}{N}\sum_{j=1}^{N}V_{ij},\quad 1\leq i\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|\mathcal{A}\right|},$ (27)
$S_{i}=\sqrt{\frac{1}{N-1}\sum_{j=1}^{N}\left(V_{ij}-\overline{V}_{i}\right)^{2}},\quad 1\leq i\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left|\mathcal{A}\right|},$ (28)
@@
$\mathcal{D}_{1}=\bigl{\{}\bigl{(}\mathbf{x}_{i},\overline{V}_{i}\bigr{)}\bigr{\}}_{1\leq i\leq\left|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{A}}\right|},\quad\mathcal{D}_{2}=\bigl{\{}\bigl{(}\mathbf{x}_{i},\overline{V}_{i}/S_{i}\bigr{)}\bigr{\}}_{1\leq i\leq\left|{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{A}}\right|}.$ (29)
@@ 4.3 Solution & Analysis
[…]
@@
$\mathbf{x}\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{X}}.$ (31)
@@ 5.1.1 Scenario Set-up
Risk free rate $r_{f}$ $\ln\left(1+0.03\right)$
@@
The parameters for the MER—$\alpha_{\text{str}}$, $\alpha_{\text{adcs}}$, and $\kappa$—are taken from Saleh et al. [7]. Additionally, $a_{\text{prop}}$ and $b_{\text{prop}}$ are determined by fitting the model presented in Eq. (6) to the mass data of high Earth orbit satellites [30]. The reference masses of payload ($m_{\text{pl},\text{ref}}$) and other subsystems ($m_{\text{base},\text{ref}}$) are set to 500 kg and 600 kg, respectively. With a design lifetime of $T_{\text{life}}=15$ years and a design propellant mass ($m_{p,\text{des}}$) 3,500 kg, the resulting dry mass ($m_{\text{dry}}$) is 1,930 kg, which is within a reasonable range. The parameters for the reliability function—$\alpha_{\text{rel}}$, $\beta_{\text{rel},1}$, $\beta_{\text{rel},2}$, $\theta_{\text{rel},1}$, and $\theta_{\text{rel},2}$—are adopted from Saleh and Castet [33]. Finally, the launch cost is set within a reasonable range, referring to current and near-future vehicles capable of delivering a GEO communication satellite to GTO—Falcon 9, Ariane 6, and Atlas V [34, 39, 40, 41, 42], and the remaining parameters are set based on data from existing articles [8, 22, 30] to ensure they fall within a reasonable range.
@@ 5.2 GEO Communication Satellite with Chemical Propulsion: Parametric Study
We repeat the procedure performed in the baseline scenario for every combination of service capacity and cost. Capacity is scaled from the baseline level (index 1) to 4 (index 2), 7 (index 3), and 10 (index 4). Service cost is scaled from the baseline (index 1) down to 0.8, 0.6, 0.4, and 0.2 (indices 2–5). Pairing each of the four capacity levels with each of the five cost levels yields a total of 20 scenarios.
@@ 5.3 GEO Communication Satellite with Electric Propulsion: Parametric Study
Risk free rate $r_{f}$ $\ln\left(1+0.03\right)$

### v3

@@ Abstract
RelRel  = reliability function of a satellite
@@
$t$  = time in discretized frame ($T/\Delta T$)
@@ 2 Problem Description
$\underset{\mathbf{x}}{\text{maximize}}\ \mathbf{J}=\left[E_{V},\ E_{V}/\sigma_{V}\right]{\color[rgb]{0,0,0},}$ (1)
@@
$\mathbf{x}=\left[T_{\text{life}},\ m_{p,\text{des}}\right]\in{\color[rgb]{0,0,0}\mathcal{X}}.$ (2)
@@ 3.1 Satellite Sizing and Cost Elements
Here, Eqs. (5) and (6) adopt the relations referred to [7]. In particular, Eq. (5) implies $\kappa$% increase in mass for each additional year of design lifetime beyond three years. We then model $m_{\text{adcs}}$ and $m_{\text{str}}$ as proportional to the dry mass, as shown in Eq. (7).
@@ 3.2.3 Decision
Decision 2 is triggered after OOR Service is selected in Decision 1. It occurs $t_{\text{oor}}$ ($t_{\text{rep}}$) time steps before the projected depletion. This event determines the quantity of propellant for refueling. The life-extension period provided by OOR Service is determined as the period with the highest utility, similar to Decision 1. When a specific life-extension duration is selected, OOR Service with the corresponding propellant amount ($m_{p,\text{oor}}$) is planned after $t_{\text{oor}}$ time steps.
@@ 3.2.4 Action
where $\Delta V_{\text{ot}}$ and $\Delta V_{\text{ot},\text{ideal}}$ represent the actual and ideal $\Delta V$ required for orbital transfer from the initial orbit to the target orbit, respectively, and $\Delta V_{\text{err}}$ is the additional $\Delta V$ needed for injection error correction. Here, $\mathcal{H}\left(\sigma_{\text{oi}}\right)$ denotes a half-normal distribution derived by taking the absolute value of a normal distribution with mean 0 and standard deviation $\sigma_{\text{oi}}$.
@@ 4.2 Surrogate Model Formulation
$\overline{V}_{i}=\frac{1}{N}\sum_{j=1}^{N}V_{ij},\quad 1\leq i\leq{\color[rgb]{0,0,0}\left|\mathcal{A}\right|},$ (27)
$S_{i}=\sqrt{\frac{1}{N-1}\sum_{j=1}^{N}\left(V_{ij}-\overline{V}_{i}\right)^{2}},\quad 1\leq i\leq{\color[rgb]{0,0,0}\left|\mathcal{A}\right|},$ (28)
@@
$\mathcal{D}_{1}=\bigl\{\bigl(\mathbf{x}_{i},\overline{V}_{i}\bigr)\bigr\}_{1\leq i\leq\left|{\color[rgb]{0,0,0}\mathcal{A}}\right|},\quad\mathcal{D}_{2}=\bigl\{\bigl(\mathbf{x}_{i},\overline{V}_{i}/S_{i}\bigr)\bigr\}_{1\leq i\leq\left|{\color[rgb]{0,0,0}\mathcal{A}}\right|}.$ (29)
@@ 4.3 Solution & Analysis
$\underset{\mathbf{x}}{\text{maximize}}\ \mathbf{\hat{J}}\left(\mathbf{x}\right)=\left[\hat{J}_{1},\hat{J}_{2}\right]{\color[rgb]{0,0,0},}$ (30)
@@
$\mathbf{x}\in{\color[rgb]{0,0,0}\mathcal{X}}.$ (31)
@@ 5.1.1 Scenario Set-up
Risk free rate $r_{f}$ $\ln\left(1+0.03\right)$ -
@@
[…]
@@ 5.2 GEO Communication Satellite with Chemical Propulsion: Parametric Study
We repeat the procedure performed in the baseline scenario for every combination of service capacity and cost as presented in Table 7. Capacity is scaled from the baseline level (index 1) to 4 (index 2), 7 (index 3), and 10 (index 4). Service cost is scaled from the baseline (index 1) down to 0.8, 0.6, 0.4, and 0.2 (indices 2–5). Pairing each of the four capacity levels with each of the five cost levels yields a total of 20 scenarios.
@@ 5.3 GEO Communication Satellite with Electric Propulsion: Parametric Study
Risk free rate $r_{f}$ $\ln\left(1+0.03\right)$ -